Hi,
I need help with the following three fraction calculations.
1. 12 1/2 + 1/ 1/3
The confusing part here for me is, the 1/2 under the 1, how do I make the denominators the same?
2. 2/ 1/2 + 1/1/5
3. 3/7 x 2 x square root 49 divided 1/9 ?
(sorry I dont know how to make the root or divided sign)
Thanks

Answer by stanbon(75887)

1. 12 1/2 + 1/ 1/3
-------------
= (2*12+1)/2 + 1*(3/1)
---
= (25/2) + 3
---
= (25/2) + (6/2)
= 31/2
----------------------------

The confusing part here for me is, the 1/2 under the 1, how do I make the denominators the same?
2. 2/ 1/2 + 1/1/5
Invert the denominators and multiply:
= (2*2) + (1*5)
= 9
==============================

3. 3/7 x 2 x square root 49 divided 1/9 ?
---
(3/7)*2*sqrt(49)/(1/9)
---------------------------
= (6/7)*(7*9)
= (6/9)
= 2/3
===================================
Cheers,
Stan H.